# Moduli stacks of Higgs bundles on stable curves

### v2

@@ 7.1.1 Versal deformations of the stabalization morphism
  1. 1.
    there exists a versal deformation space of the nodal curve $\mathcal{X}$ which is isomorphic to ${\rm Spec}\,\mathbb{k}[|z_{1},\dots,z_{l}|][|z_{l+1},\dots,z_{N}|]$, where $N:=3g-3$ (here $g$ is the genus of the curves) and $z_{i}$ is the equation of the $i$-th node of $\mathcal{X}$ for $i=1,\dots,l$, and
  2. 2.
    there exits a versal deformation space of the nodal curve $\mathfrak{X}$ which is isomorphic to ${\rm Spec}\,\mathbb{k}[|\{\{z_{ij}\}^{l}_{i=1}\}^{\iota_{i}}_{j=1}|][|z_{l+1},\dots,z_{N}|]$, where $z_{ij}=0$ is the equation of the node $d_{ij}$ for $i=1,\dots,l$ and $j=1,\dots,\iota_{i}$. Moreover, there is a morphism between the versal deformation spaces, which given is as follows.
@@
    $z_{i}\mapsto z_{i1}\cdots z_{i\iota_{i}}\hskip 30.00005pt\forall i=1,\dots,l,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}$
@@
[…]
@@
is a versal deformation family of curves for a node (see [14, Theorem 14.1]). If we consider the basic log structure package on ${\rm Spec}\,(\mathbb{k}[x,y,a]/(xy-a))/\operatorname{Spec}\leavevmode\nobreak\ (\mathbb{k}[a])$, the basic log structures on $\operatorname{Spec}\leavevmode\nobreak\ \mathbb{k}[a]$, agrees by Kato [21, Lemma 4.4] with the divisorial log structure given by the boundary divisor (i.e., the vanishing locus of the element $a\in\mathbb{k}[a]$). Now from [21, Lemma 2.1, Lemma 2.2, Proposition 2.3], we can find an étale neighbourhood $V$ of the point $p\in\overline{\mathcal{M}_{g}}$, and define a log structure by taking the amalgamated sum of the log structures contributed by each of the nodes in $\{q_{1},\dots,q_{n}\}$. But this is the same as the log structures induced by the pullback of the boundary divisor in $\overline{\mathcal{M}_{g}}$. This completes the argument. Similarly, the two log structures on $\mathcal{M}^{ss}_{g}$ are the same.
@@
Notice that the rightmost vertical morphism is a morphism of log schemes. Since the right commutative diagram is commutative and the log structures on $V$ and $W$ are isomorphic to the pullback of the log structures from $\operatorname{Spec}\leavevmode\nobreak\ \mathbb{k}[\{\{z_{ij}\}^{l}_{i=1}\}^{\iota_{i}}_{j=1}][\{z_{k}\}^{N}_{k=l+1}]$ and $\mathbb{k}[z_{1},\dots,z_{l}][z_{l+1},\dots,z_{N}]$, therefore we have the structure of a log-morphism on $V\rightarrow W$.
@@
[…]
where $e_{i}$ is sent to $z_{i}$ and $e_{ij}$ is sent to $z_{ij}$, where $i=1,\dots,l$. ∎
7.2 Relative log-cotangent complex of the map ${\mathcal{M}^{ss}_{g}}\stackrel{{\scriptstyle f}}{{\longrightarrow}}\overline{\mathcal{M}_{g}}$
@@
${\mathcal{D}_{g}}$${\mathcal{C}_{g}}$${\mathcal{M}^{ss}_{g}}$${\overline{\mathcal{M}_{g}}}$$\scriptstyle{\tilde{p}}$$\scriptstyle{\tilde{f}}$$\scriptstyle{p}$$\scriptstyle{f}$ (7.7)
These families of semi-stable curves as discussed in Remark 7.1 have relative log structure packages, defining log structures on all the spaces in this diagram, for which the projection morphisms $p$ and $\tilde{p}$ are log-smooth. The log-tangent complex of $\overline{\mathcal{M}_{g}}$ is isomorphic to $Rp_{*}(\omega^{\vee}_{\mathcal{C}_{g}/\overline{\mathcal{M}_{g}}})[1]$ and the log-tangent complex of $\mathcal{M}^{ss}_{g}$ is isomorphic to $R\tilde{p}_{*}(\omega^{\vee}_{\mathcal{D}_{g}/{\mathcal{M}^{ss}_{g}}})[1]$. Now notice that $\tilde{f}^{*}\omega^{\vee}_{\mathcal{C}_{g}/\overline{\mathcal{M}_{g}}}\cong\omega^{\vee}_{\mathcal{D}_{g}/{\mathcal{M}^{ss}_{g}}}$ and that the fibers of $\tilde{f}$ are connected and so $R\tilde{f}_{*}\mathcal{O}\cong\mathcal{O}$. Hence, $R\tilde{f}_{*}\omega^{\vee}_{\mathcal{C}_{g}/\overline{\mathcal{M}_{g}}}\cong\omega^{\vee}_{\mathcal{D}_{g}/{\mathcal{M}^{ss}_{g}}}$. Therefore we have,
@@
$f^{*}\mathbb{L}^{\log}_{\overline{\mathcal{M}_{g}}}\longrightarrow\mathbb{L}^{\log}_{{\mathcal{M}^{ss}_{g}}}\longrightarrow\mathbb{L}^{\log}_{f}\longrightarrow f^{*}\mathbb{L}^{\log}_{\overline{\mathcal{M}_{g}}}[1]$ (7.8)
@@
7.3 Relative logarithmic Dolbeault shape and shifted symplectic forms
@@
There is a 00-shifted relative log-symplectic form on $\mathcal{M}^{Dol}_{g}$ (relative to the moduli stack of stable curves $\overline{\mathcal{M}_{g}}$).
@@
[…]
$\pi^{*}\mathbb{L}^{\log}_{\mathcal{M}^{ss}_{g}/\overline{\mathcal{M}_{g}}}\longrightarrow\mathbb{L}^{\log}_{\mathcal{M}^{Dol}_{g}/\overline{\mathcal{M}_{g}}}\longrightarrow\mathbb{L}^{\log}_{\mathcal{M}^{Dol}_{g}/\overline{\mathcal{M}_{g}}}\longrightarrow\pi^{*}\mathbb{L}^{\log}_{\mathcal{M}^{ss}_{g}/\overline{\mathcal{M}_{g}}}[1].$ (7.9)
But since $\mathbb{L}^{\log}_{\mathcal{M}^{ss}_{g}/\overline{\mathcal{M}_{g}}}\cong 0$ (Proposition 7.4), therefore the relative log cotangent complex of the composite morphism is isomorphic to the relative log cotangent complex of the morphism $\pi$. But notice that the log structure of $\mathcal{M}^{Dol}_{g}$ is pulled back of the log structure of $\mathcal{M}^{ss}_{g}$ via the map $\pi$. Therefore the relative log-cotangent complex of the morphism $\pi$ is isomorphic to the relative cotangent complex of the morphism $\pi$. Now since the relative logarithmic Dolbeault stack over the moduli stacks $\mathcal{M}^{Dol}_{g}$ is $\mathcal{O}$-compact and $\mathcal{O}$-oriented (Theorem 4.7); therefore $\mathcal{M}^{Dol}_{g}$ has a 00-shifted relative symplectic form over $\mathcal{M}^{ss}_{g}$, which is a 00-shifted relative log-symplectic form viewed over the moduli stack of stable curves $\overline{\mathcal{M}_{g}}$. ∎
8 Appendix: Classical Artin stack of Gieseker-Higgs bundles and its local properties
In this section, we will define and construct the relative classical Artin stack of Gieseker-Higgs bundles and study its local properties. The main results of the appendix are Lemma 8.10 and Lemma 8.11. In the first lemma, we prove that the stack of Gieseker vector bundles is an almost very good stack ([37, Definition 2.1.2]). We use this in Lemma 8.11 to show that the classical stack of Gieseker-Higgs bundles is an irreducible, local complete intersection.
8.1 The stacks of Gieseker vector bundles and Gieseker Higgs bundles
Definition 8.1.
(Gieseker vector bundle) A vector bundle $\mathcal{E}$ of rank $n$ on $X_{r}$ with $r\geq 1$ is called a Gieseker vector bundle if
  1. 1.
@@ 8.1 The stacks of Gieseker vector bundles and Gieseker Higgs bundles
Any vector bundle on $X_{0}$ is called a Gieseker vector bundle. In the literature, a Gieseker vector bundle is also called an admissible vector bundle.
A Gieseker vector bundle $(X_{r},\mathcal{E})$ is called a stable Gieseker vector bundle if $\pi_{r*}\mathcal{E}$ is a stable torsion-free sheaf on the irreducible nodal curve $X_{0}$, where $\pi_{r}:X_{r}\longrightarrow X_{0}$ is the natural contraction map.
A (stable) Gieseker vector bundle on a modification $\mathfrak{X}_{T}$ (see Definition 3.2) is a vector bundle such that its restriction to each $(\mathfrak{X}_{T})_{t}$ is a (stable) Gieseker vector bundle.
Remark 8.2.
@@
Definition 8.3.
(Gieseker-Higgs bundle) A Gieseker-Higgs bundle on $\mathfrak{X}_{T}$ is a pair $(\mathcal{E}_{T},\phi_{T})$, where $\mathcal{E}_{T}$ is a vector bundle on $\mathfrak{X}_{T}$, and $\phi_{T}:\mathcal{E}_{T}\longrightarrow\mathcal{E}_{T}\otimes\omega_{\mathfrak{X}_{T}/T}$ is an $\mathcal{O}_{\mathfrak{X}_{T}}$module homomorphism satisfying the following
  1. 1.
@@
8.1.1 Stack of torsion-free Hitchin pairs
@@
$TFH(\mathcal{X}/S):{Sch}/S\rightarrow{Groupoids}$
@@
It is an Artin stack. For the construction of an atlas for $TFH(\mathcal{X}/S)$, we refer to [1, 5.0.7. The total family construction.]. Following the notation from [1], we denote it by $\coprod_{m\geq m_{0}}R^{\Lambda,m}_{S}$. The superscript “$\Lambda$” is because for the construction of the quot scheme of torsion-free Hitchin pairs one views them as $\Lambda$-modules, where $\Lambda=\operatorname{Sym}\leavevmode\nobreak\ (\omega^{\vee}_{\mathcal{X}/S})$ [36]. We denote the stack of families of torsion-free sheaves over $\mathcal{X}/S$ of rank $n$ and degree $d$ by $TF(\mathcal{X}/S)$. The stack $TF(\mathcal{X}/S)$ is a reduced and irreducible Artin stack.
8.1.2 Ordinary Artin Stack of Gieseker-Higgs bundles
Given any derived Artin stack $F:\textsf{cdga}\longrightarrow\mathbb{S}$ one can define a classical/ordinary Artin stack by considering the composition functor $F^{cl}:\textsf{alg}_{\mathbb{k}}\longrightarrow\textsf{cdga}_{\mathbb{k}}\longrightarrow\mathbb{S}$. We call $F^{cl}$ as the underlying classical Artin stack of $F$. Following this notation, we denote the underlying classical Artin stack of the derived stack $M$ of Higgs bundles over the family of curves $\mathcal{X}_{\mathfrak{M}}/\mathfrak{M}$ by $M^{cl}$.
Let us denote by $\operatorname{Coh}(\mathcal{X}/S)$ the Artin stack of coherent $\mathcal{O}_{\mathcal{X}}-$ modules which are flat over $S$. There is a natural map
$\theta:M^{cl}\longrightarrow\operatorname{Coh}(\mathcal{X}/S)$
which is given by the push forward of the underlying bundle $\pi_{*}\mathcal{E}$, where $\pi:\mathcal{X}^{mod}\longrightarrow\mathcal{X}$ is the modification morphism.
Consider the open sub-stack $TF(\mathcal{X}/S)\subset\operatorname{Coh}(\mathcal{X}/S)$ consisting of torsion-free sheaves. The stack $\theta^{-1}(TF(\mathcal{X}/S))$ is an open substack of $M^{cl}$. Let us consider the natural map $\pi^{*}\pi_{*}\mathcal{E}\longrightarrow\mathcal{E}$ over the universal curve $\mathfrak{X}$ over $M^{cl}$. Consider the sub-stack of $M^{cl}$ where the map is surjective. It is again an open sub-stack. Let us denote it by $M^{gg}$ (the superscript ”gg” stands for globally generated). Then we define an open sub-stack

### v3

@@ 7.1.1. Versal deformations of the stabalization morphism
  1. (1)
    there exists a versal deformation space of the nodal curve $\mathcal{X}$ which is isomorphic to ${\rm Spec}\,\mathbb{k}[|z_{1},\dots,z_{l}|][|z_{l+1},\dots,z_{N}|]$, where N:=3g3N:=3g-3 (here $g$ is the genus of the curves) and $z_{i}$ is the equation of the $i$-th node of $\mathcal{X}$ for $i=1,\dots,l$, and
  2. (2)
    there exits a versal deformation space of the nodal curve $\mathfrak{X}$ which is isomorphic to ${\rm Spec}\,\mathbb{k}[|\{\{z_{ij}\}^{l}_{i=1}\}^{\iota_{i}}_{j=1}|][|z_{l+1},\dots,z_{N}|]$, where $z_{ij}=0$ is the equation of the node $d_{ij}$ for $i=1,\dots,l$ and $j=1,\dots,\iota_{i}$. Moreover, there is a morphism between the versal deformation spaces, which is given as follows.
@@
    $z_{i}\mapsto z_{i1}\cdots z_{i\iota_{i}}\hskip 30.00005pt\forall i=1,\dots,l,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \text{and}$
@@
    This morphism is the local analytic picture of the stabilisation morphism $\pi:\mathcal{M}^{ss}_{g}\longrightarrow\overline{\mathcal{M}_{g}}$ at a given point of $\mathcal{M}^{ss}_{g}$.
@@
(7.3) $\mathcal{C}_{g}$$U_{i}$${\rm Spec}\,(B[x,y]/(xy-b_{i}))$${\rm Spec}\,(\mathbb{k}[x,y,a]/(xy-a))$$\overline{\mathcal{M}_{g}}$${\rm Spec}\,B$${\rm Spec}\,\mathbb{k}[a]$$\scriptstyle{etale}$$\scriptstyle{etale}$$\scriptstyle{etale}$
@@
is a versal deformation family of curves for a node (see [14, Theorem 14.1]). If we consider the basic log structure package on ${\rm Spec}\,(\mathbb{k}[x,y,a]/(xy-a))/\operatorname{Spec}\penalty 10000\ (\mathbb{k}[a])$, the basic log structures on $\operatorname{Spec}\penalty 10000\ \mathbb{k}[a]$, agrees by Kato [15, Lemma 4.4] with the divisorial log structure given by the boundary divisor (i.e., the vanishing locus of the element $a\in\mathbb{k}[a]$). Now from [15, Lemma 2.1, Lemma 2.2, Proposition 2.3], we can find an étale neighbourhood $V$ of the point $p\in\overline{\mathcal{M}_{g}}$, and define a log structure by taking the amalgamated sum of the log structures contributed by each of the nodes in $\{q_{1},\dots,q_{n}\}$. But this is the same as the log structures induced by the pullback of the boundary divisor in $\overline{\mathcal{M}_{g}}$. This completes the argument. Similarly, the two log structures on $\mathcal{M}^{ss}_{g}$ are the same.
@@
Notice that the rightmost vertical morphism is a morphism of log schemes. Since the right commutative diagram is commutative and the log structures on $V$ and $W$ are isomorphic to the pullback of the log structures from $\operatorname{Spec}\penalty 10000\ \mathbb{k}[\{\{z_{ij}\}^{l}_{i=1}\}^{\iota_{i}}_{j=1}][\{z_{k}\}^{N}_{k=l+1}]$ and $\mathbb{k}[z_{1},\dots,z_{l}][z_{l+1},\dots,z_{N}]$, therefore we have the structure of a log-morphism on $V\rightarrow W$.
@@
(7.6) $e_{i}$$\sum^{\iota_{i}}_{j=1}e_{ij}$$\mathbb{N}^{\oplus l}$$\oplus^{l}_{i=1}(\mathbb{N}^{\oplus\iota_{i}})$$\mathbb{k}[|z_{1},\dots,z_{l}|][|z_{l+1},\dots,z_{N}|]$$\mathbb{k}[|\{\{z_{ij}\}^{l}_{i=1}\}^{\iota_{i}}_{j=1}|][|\{z_{k}\}^{N}_{k=l+1}|]$$scriptstyle{z_{i}=z_{i1}\cdots z_{i\iota_{i}}}$
where $e_{i}$ is sent to $z_{i}$ and $e_{ij}$ is sent to $z_{ij}$, where $i=1,\dots,l$.
The above map can be seen a tensor products of the following maps over the field $k$ along with an extra tensor product with $k[|z_{l+1},\dots,z_{N}|]$, where $k[|z_{l+1},\dots,z_{N}|]$ does not contribute at all in the log structure.
(7.7) $e_{i}$$\sum^{\iota_{i}}_{j=1}e_{ij}$$\mathbb{N}$$\mathbb{N}^{\oplus\iota_{i}}$$\mathbb{k}[|z_{i}|]$$\mathbb{k}[|\{z_{ij}\}^{\iota_{i}}_{j=1}|]$$scriptstyle{z_{i}=z_{i1}\cdots z_{i\iota_{i}}}$
where $e_{i}$ is sent to $z_{i}$ and $e_{ij}$ is sent to $z_{ij}$.
Now it is straightforward to check that each of these maps is log-smooth using Kato’s criterion (c.f. [16, Theorem 3.5]). Since the product of log-smooth maps is log-smooth, it follows that the map $\mathcal{M}^{ss}_{g}\longrightarrow\overline{\mathcal{M}_{g}}$ is log-smooth. ∎
7.2. Relative log-cotangent complex of the map ${\mathcal{M}^{ss}_{g}}\stackrel{{\scriptstyle f}}{{\longrightarrow}}\overline{\mathcal{M}_{g}}$
@@
(7.8) $\mathcal{D}_{g}$$\mathcal{C}_{g}$$\mathcal{M}^{ss}_{g}$$\overline{\mathcal{M}_{g}}$$\scriptstyle{\tilde{p}}$$\scriptstyle{\tilde{f}}$$\scriptstyle{p}$$\scriptstyle{f}$
These families of semi-stable curves as discussed in Remark 7.1 have relative log structure packages, defining log structures on all the spaces in this diagram, for which the projection morphisms $p$ and $\tilde{p}$ are log-smooth ([15, Theorem 2.1., and “Gobal Construction”, Page 227]). The log-tangent complex of $\overline{\mathcal{M}_{g}}$ is isomorphic to $Rp_{*}(\omega^{\vee}_{\mathcal{C}_{g}/\overline{\mathcal{M}_{g}}})[1]$ and the log-tangent complex of $\mathcal{M}^{ss}_{g}$ is isomorphic to $R\tilde{p}_{*}(\omega^{\vee}_{\mathcal{D}_{g}/{\mathcal{M}^{ss}_{g}}})[1]$. Now notice that $\tilde{f}^{*}\omega^{\vee}_{\mathcal{C}_{g}/\overline{\mathcal{M}_{g}}}\cong\omega^{\vee}_{\mathcal{D}_{g}/{\mathcal{M}^{ss}_{g}}}$ (see [19, Page 169]) and that the fibers of $\tilde{f}$ are connected and so $R\tilde{f}_{*}\mathcal{O}\cong\mathcal{O}$. Hence, $R\tilde{f}_{*}\omega^{\vee}_{\mathcal{C}_{g}/\overline{\mathcal{M}_{g}}}\cong\omega^{\vee}_{\mathcal{D}_{g}/{\mathcal{M}^{ss}_{g}}}$. Therefore we have,
@@
(7.9) $f^{*}\mathbb{L}^{\log}_{\overline{\mathcal{M}_{g}}}\longrightarrow\mathbb{L}^{\log}_{{\mathcal{M}^{ss}_{g}}}\longrightarrow\mathbb{L}^{\log}_{f}\longrightarrow f^{*}\mathbb{L}^{\log}_{\overline{\mathcal{M}_{g}}}[1]$
@@
7.3. Relative logarithmic Dolbeault shape and shifted symplectic forms
@@
There is a 0-shifted relative log-symplectic form on $\mathcal{M}^{Dol}_{g}$ (relative to the moduli stack of stable curves $\overline{\mathcal{M}_{g}}$).
@@
Consider the composite morphism $\mathcal{M}^{Dol}_{g}\xrightarrow{\pi}\mathcal{M}^{ss}_{g}\xrightarrow{f}\overline{\mathcal{M}_{g}}$. We have a distinguished triangle
(7.10) $\pi^{*}\mathbb{L}^{\log}_{\mathcal{M}^{ss}_{g}/\overline{\mathcal{M}_{g}}}\longrightarrow\mathbb{L}^{\log}_{\mathcal{M}^{Dol}_{g}/\overline{\mathcal{M}_{g}}}\longrightarrow\mathbb{L}^{\log}_{\mathcal{M}^{Dol}_{g}/\overline{\mathcal{M}_{g}}}\longrightarrow\pi^{*}\mathbb{L}^{\log}_{\mathcal{M}^{ss}_{g}/\overline{\mathcal{M}_{g}}}[1].$
But since $\mathbb{L}^{\log}_{\mathcal{M}^{ss}_{g}/\overline{\mathcal{M}_{g}}}\cong 0$ (Proposition 7.4), therefore the relative log cotangent complex of the composite morphism is isomorphic to the relative log cotangent complex of the morphism $\pi$. But notice that the log structure of $\mathcal{M}^{Dol}_{g}$ is pulled back of the log structure of $\mathcal{M}^{ss}_{g}$ via the map $\pi$. Therefore the relative log-cotangent complex of the morphism $\pi$ is isomorphic to the relative cotangent complex of the morphism $\pi$. Now since the relative logarithmic Dolbeault stack over the moduli stacks $\mathcal{M}^{Dol}_{g}$ is $\mathcal{O}$-compact and $\mathcal{O}$-oriented (Theorem 4.7); therefore $\mathcal{M}^{Dol}_{g}$ has a 0-shifted relative symplectic form over $\mathcal{M}^{ss}_{g}$, which is a 0-shifted relative log-symplectic form viewed over the moduli stack of stable curves $\overline{\mathcal{M}_{g}}$. ∎
Appendix A Classical Artin stack of Gieseker-Higgs bundles and its local properties
In this appendix we study the classical truncation $M^{cl}_{Gie}$ of the derived moduli stack of Higgs bundles (Section 3), restricted to the Gieseker locus (Definition A.3). Our main results are:
- The stack of Gieseker vector bundles $N_{Gie}$ (and in particular its closed fiber $N^{cl}_{Gie,0}$) is an almost very good stack in the sense of Soibelman [39, Definition 2.1.2].
- The classical stack of Gieseker-Higgs bundles over the special fiber, $M^{cl}_{Gie,0}$, is an irreducible local complete intersection of pure dimension $2\dim N^{cl}_{Gie,0}+1$.
These facts are used in Section 5 to deduce flatness of the Hitchin map on the Gieseker locus.
A.1. The classical Artin stacks of Gieseker vector bundles and Gieseker Higgs bundles
We now describe the underlying classical Artin stacks associated to the derived moduli problems considered earlier. Let $N_{Gie}$ denote the classical Artin stack parameterizing Gieseker vector bundles, and let $M^{cl}_{Gie}$ denote the classical Artin stack parameterizing Gieseker–Higgs bundles (both defined over the base $S$).
A.1.1. Gieseker vector bundles
Definition A.1.
A vector bundle $\mathcal{E}$ of rank $n$ on $X_{r}$ with $r\geq 1$ is called a Gieseker vector bundle if
  1. (1)
@@ A.1.1. Gieseker vector bundles
Any vector bundle on $X_{0}$ is called a Gieseker vector bundle. In the literature, a Gieseker vector bundle is also called an admissible vector bundle. A Gieseker vector bundle $(X_{r},\mathcal{E})$ is called a stable Gieseker vector bundle if $\pi_{r*}\mathcal{E}$ is a stable torsion-free sheaf on the irreducible nodal curve $X_{0}$, where $\pi_{r}:X_{r}\longrightarrow X_{0}$ is the natural contraction map. A (stable) Gieseker vector bundle on a modification $\mathfrak{X}_{T}$ (see Definition 3.3) is a vector bundle such that its restriction to each $(\mathfrak{X}_{T})_{t}$ is a (stable) Gieseker vector bundle.
Remark A.2.
@@
Definition A.3.
A Gieseker-Higgs bundle on $\mathfrak{X}_{T}$ is a pair $(\mathcal{E}_{T},\phi_{T})$, where $\mathcal{E}_{T}$ is a vector bundle on $\mathfrak{X}_{T}$, and $\phi_{T}:\mathcal{E}_{T}\longrightarrow\mathcal{E}_{T}\otimes\omega_{\mathfrak{X}_{T}/T}$ is an $\mathcal{O}_{\mathfrak{X}_{T}}$ -module homomorphism satisfying the following
  1. (1)
@@
A.1.2. Stack of torsion-free Hitchin pairs
@@
$\mathsf{TFH}(\mathcal{X}/S):{Sch}/S\rightarrow{Groupoids}$
@@
It is an Artin stack. For the construction of an atlas for $\mathsf{TFH}(\mathcal{X}/S)$, we refer to [2, 5.0.7. The total family construction.]. Following the notation from [2], we denote it by $\coprod_{m\geq m_{0}}R^{\Lambda,m}_{S}$. The superscript $\Lambda$” is because for the construction of the quot scheme of torsion-free Hitchin pairs one views them as $\Lambda$-modules, where $\Lambda=\operatorname{Sym}\penalty 10000\ (\omega^{\vee}_{\mathcal{X}/S})$ [36]. We denote the stack of families of torsion-free sheaves over $\mathcal{X}/S$ of rank $n$ and degree $d$ by $TF(\mathcal{X}/S)$. The stack $TF(\mathcal{X}/S)$ is a reduced and irreducible Artin stack.
A.1.3. Classical Artin stack of Gieseker-Higgs bundles
Given any derived Artin stack $F:\textsf{cdga}\longrightarrow\mathbb{S}$ one can define a classical/ordinary Artin stack by considering the composition functor $F^{cl}:\textsf{alg}_{\mathbb{k}}\longrightarrow\textsf{cdga}_{\mathbb{k}}\longrightarrow\mathbb{S}$. We call $F^{cl}$ as the underlying classical Artin stack of $F$. Following this notation, we denote the underlying classical Artin stack of the derived stack $M$ of Higgs bundles over the family of curves $\mathcal{X}_{\mathfrak{M}}/\mathfrak{M}$ by $M^{cl}$. Let us denote by $\operatorname{\mathsf{Coh}}(\mathcal{X}/S)$ the Artin stack of coherent $\mathcal{O}_{\mathcal{X}}$- modules that are flat over $S$. There is a natural map
$\theta:M^{cl}\longrightarrow\operatorname{\mathsf{Coh}}(\mathcal{X}/S)$
which is given by the push forward of the underlying bundle $\pi_{*}\mathcal{E}$, where $\pi:\mathfrak{X}\longrightarrow\mathcal{X}$ is the modification morphism. Consider the open sub-stack $TF(\mathcal{X}/S)\subset\operatorname{\mathsf{Coh}}(\mathcal{X}/S)$ consisting of torsion-free sheaves. The stack $\theta^{-1}(TF(\mathcal{X}/S))$ is an open substack of $M^{cl}$. Consider the natural map $\pi^{*}\pi_{*}\mathcal{E}\longrightarrow\mathcal{E}$ on the universal curve $\mathfrak{X}$ on $M^{cl}$. Consider the sub-stack of $M^{cl}$ where the map is surjective. It is again an open sub-stack. Let us denote it by $M^{gg}$ (the superscript ”gg” stands for globally generated). Then we define an open sub-stack